Jeorge is 3 years older than his sister Janet.In 7 years she will be six-seven of his age.How old are they?
Answer by mananth(16949)

sister =x
Jeorge = x+3
..
7 years hence
sister = x+7
jeorge = x+3+7=x+10
..
x+7 = 6(x+10)/7
7(x+7)=6(x+10)
7x+49=6x+60
7x-6x=60-49
x=11 sister's age
..s
Jeorge 11+3 =14 year